{x}'' - {x}' - 10x = 100t^{3} + 12.5; x(0) = 5, {x}'(0) = -0.5

How would I find the total solution to this initial value problem?
Answer #1

x_x- 10x = 100 t3 +12.5. dx - dx - 10% = 100* +12.5 op (D?-D-10)x= 100 t3 +12.5 D = d. The auxiliary equation is, m?_m- 10= - - 1.25 1000 t3 - 300 +² +660 +-126 100 = - 10 t + 3 t² _ 6.6t+1.26-1.25 = - 10 it²+ 3+² – 6.6 t + 0.01 So, the complete s